# Foundation Model for Polycrystalline Material Informatics

### v2

@@ 2.3.3 Downstream task II: Nonlinear homogenized response prediction
Beyond homogenized stiffness prediction, the second downstream objective extends the pretrained encoder toward nonlinear homogenized response prediction by predicting the parameters of the ODMN [Wei01]. The predicted ODMN parameters are subsequently used to evaluate nonlinear stress–strain responses. This downstream task assesses whether the learned latent representations can be transferred from linear elastic property prediction to the more challenging task of surrogate model parameter prediction.
@@ 4 Conclusions
This study introduced a three-dimensional polycrystal foundation model pretrained on a large corpus of FCC microstructures whose crystallographic orientations span the texture hull. Through a self-supervised masked reconstruction strategy with masking ratios ranging from 20% to 90%, the encoder was trained on 100,000100{,}000 voxel-based microstructures and learned a physically meaningful latent representation. The resulting latent distribution exhibits an approximately uniform circular topology, indicating that the pretrained encoder provides a smooth and isotropic embedding of the entire pretraining dataset.

### v3

@@ 2.3.3 Downstream task II: Nonlinear homogenized response prediction
Building upon the framework introduced in the previous subsection, this downstream task addresses the limited transferability of ODMN parameters by coupling the ODMN with the pretrained foundation model. Specifically, the foundation model is employed to enable systematic inference of ODMN parameters for previously unseen polycrystalline microstructures, thereby extending the pretrained encoder from linear elastic property prediction to nonlinear homogenized response prediction [Wei01]. The inferred ODMN parameters are subsequently used to evaluate nonlinear stress–strain responses under crystal plasticity. This downstream task, therefore, assesses whether the learned latent representations can be transferred from elastic property prediction to the more challenging task of surrogate model parameter inference.
@@ 4 Conclusions
This study introduced a three-dimensional polycrystal foundation model pretrained on a large corpus of FCC microstructures whose crystallographic orientations span the texture hull. Through a self-supervised masked reconstruction strategy with masking ratios ranging from 20% to 90%, the encoder was trained on 100,000100{,}000 voxel-based microstructures and learned a physically meaningful latent representation. The resulting latent distribution is broadly and smoothly populated, indicating a continuous embedding of the pretraining dataset.
@@ Appendix B Generalized Hooke’s law
Declarations
The authors declare that there are no competing interests.
Data Availability
The datasets used in this study are publicly available at the Zenodo repository: https://doi.org/10.5281/zenodo.18009973.
Code Availability
The code used in this study can be accessed at the following GitHub repository: https://github.com/BerryWei/Foundation-Model-for-Polycrystalline-Material-Informatics.